A 300mm square column is reinforced with 3Φ16 bars at each of the two critical faces. If C25 concrete and steel with fyk=300MPa is used, compute the ultimate axial compressive force the section can safely resist.



1
Expert's answer
2022-03-19T02:46:38-0400

CS = F ÷ A, Where;


CS is the compressive strength,


F is the force or load at point of failure and


A is the initial cross-sectional surface area.


CS= "\frac{300MPa}{3Φ16(300mm)}"


CS= "\frac{300}{77.67}"


CS= 3.86 Pa